A small mailbag is released from a helicopter that is descending steadily at 1.50 m/s.

(a) After 3.00 s, what is the speed of the mailbag?
v = m/s

(b) How far is it below the helicopter?
d = m

(c) What are your answers to parts (a) and (b) if the helicopter is rising steadily at 1.50 m/s?

v = m/s
d = m
Answer #1

Dear student,

Find this solution, and RATE IT ,If you find it is helpful .your rating is very important to me.If any incorrectness ,kindly let me know I will rectify them soon.

Thanks for asking ..Ah=d=h-h=39.6-(-4.5)=44.1m) Vo = 105 misec a) Vj = Initial velocity of mail box= 1-5 mlec, t=3 sec, V = Vi tgt = 1.5+9.8x3 =